sin 60 cos30 cos60 sin30 sin30 o o o o o
1
Expert's answer
2020-07-28T18:39:46-0400


Let X = "\sin 60\degree * \cos 30\degree * \cos 60\degree * \sin 30\degree * \sin 30\degree"


"\sin 60\degree = \sqrt{3}/2"

"\cos 30\degree = \sqrt{3}/2"

"\cos 60\degree = 1/2"

"\sin 30\degree = 1/2"


So substituting the above values in X

"\therefore" X = "\sqrt{3}/2" * "\sqrt{3}/2" * "0.5" * "0.5" * "0.5"


= 3/32